OpenStudy (anonymous):

I need help in finding the absolute max of T(x,y) = x^2 - y^2 - 2x +1, I have the gradient, just need to know how to test the bounds.

OpenStudy (anonymous):

Bounds are between 0 \[0 \le x \le 2\]

OpenStudy (anonymous):

And \[-1 \le y \le 1\]